# Deep Pre-trained Time Series Features for Tree Species Classification in the Dutch Forest Inventory

### v1

@@ 2.3 Remote Sensing Time Series
Since mono-temporal data cannot adequately capture seasonal leaf phenology for tree species classification, we utilized time-series data spanning January through December 2020. This temporal range aligns with both the Pretrained Remote Sensing Transformer (Presto) model requirements (Tseng et al., 2024) and previous Dutch tree species classification research (Francini et al., 2024).
@@
Input Data Modalities S2, S1, ERA5, SRTM. This study employs a comprehensive multi-source remote sensing (RS) approach that integrates satellite imagery and environmental data to characterize forest vegetation across space and time. The model relies on several key data sources: Sentinel-1 SAR GRD data with VH and VV polarizations at a 6-day revisit frequency; Sentinel-2 multispectral Level 1C data with 5–10 day revisit frequency, including 10 spectral bands (B2–B8, B8A, B11, B12) and NDVI computed during preprocessing; ERA5 climate data providing monthly measurements of 2m air temperature and total precipitation; and static SRTM terrain data comprising elevation and slope.

### v2

@@ 3.3 Remote Sensing Time Series for Presto and harmonic + medoid calculation
We extracted time-series data spanning January through December 2020 through Google Earth Engine (Gorelick et al., 2017). This temporal range aligns with both the Pretrained Remote Sensing Transformer (Presto) model requirements (Tseng et al., 2024) and previous Dutch tree species classification research (Francini et al., 2024).
@@
Input Data Modalities S-2, S-1, ERA5, SRTM. This study employs a comprehensive multi-source remote sensing (RS) approach that integrates satellite imagery and environmental data to characterize forest vegetation across space and time. The model relies on several key data sources: S-1 SAR GRD data with VH and VV polarizations at a 6-day revisit frequency; S-2 multispectral Level 1C data with 5–10 day revisit frequency, including 10 spectral bands (B2–B8, B8A, B11, B12) and NDVI computed during preprocessing; ERA5 climate data providing monthly measurements of 2m air temperature and total precipitation; and static SRTM terrain data comprising elevation and slope. Using the plot-level coordinates, we extracted a single set of time-series data (or embeddings) for each plot; consequently, the COMB dataset comprises 13,790 plots and 13,790 corresponding samples. During preprocessing, satellite data with coarser spatial resolutions than 10 m (i.e., S-1, ERA5, and SRTM) were resampled to a common spatial resolution of 10 m to ensure consistency with Presto input(Tseng et al., 2024) and harmonic + medoid feature input(Francini et al., 2024).